OpenStudy (rebeccapink6398):

Convert 8.3 to a percent @appleboy2 could you help :D or anyone!

OpenStudy (appleboy2):

Sure!

OpenStudy (appleboy2):

Its's 830%.

OpenStudy (rebeccapink6398):

THX! could you also help me with this??? Which fraction is equivalent to 12%? I don't get it

OpenStudy (appleboy2):

Put the fractions here please.

OpenStudy (appleboy2):

Or I can just find out.

OpenStudy (rebeccapink6398):

?

OpenStudy (rebeccapink6398):

OH! one min!

OpenStudy (anonymous):

12% = 0.12 = 12/100 = 3/25

OpenStudy (appleboy2):

OK.

OpenStudy (appleboy2):

That's it.

OpenStudy (appleboy2):

Any other questions?

OpenStudy (rebeccapink6398):

A. 3/25 B. 1/8 C. 12/25 D. 49/25 there.. is that what ur looking for?

OpenStudy (appleboy2):

Too late now, Bella solved it.

OpenStudy (rebeccapink6398):

OOPS! lol I didn't see that

OpenStudy (appleboy2):

It's alright. Need any more help?

OpenStudy (rebeccapink6398):

lol

OpenStudy (rebeccapink6398):

yes please :D What is 5% of 78? A. 0.39 B. 3.9 C. 39 D. 390

OpenStudy (rebeccapink6398):

I think its B.?

OpenStudy (appleboy2):

Let's see.

OpenStudy (appleboy2):

You're correct!

OpenStudy (rebeccapink6398):

16% of what number is 120? A. 7.5 B. 19.2 C. 750 D. 7500 ~~ I think its D.?

OpenStudy (appleboy2):

Hold on.

OpenStudy (appleboy2):

It's 750.

OpenStudy (anonymous):

No, I think it is c if 120 is 16% then 120/16 = 1% so 7.5 = 1% now there's 100%, so we take one percent and multiply it by 100 which is 7.5 x 100 = 750 therefore, 120 is 16% of 750

OpenStudy (rebeccapink6398):

oh!

OpenStudy (appleboy2):

That's it.

OpenStudy (rebeccapink6398):

Could you guys also help me with my last 2 questions?

OpenStudy (appleboy2):

Sure!

OpenStudy (anonymous):

Of course!

OpenStudy (rebeccapink6398):

:D

OpenStudy (rebeccapink6398):

1. The regular price of a sweater was $21.50. The sweater was on sale for 18% off. How much was the discount? 2. Max bought a backpack that cost $40.00. The sales tax in his state is 8%. What is the final amount that Max has to pay for the backpack?

OpenStudy (appleboy2):

Discount was $3.87.

OpenStudy (anonymous):

1. I think it is 3.87 2. 3.20.

OpenStudy (appleboy2):

And 2 is correct, Bella.

OpenStudy (anonymous):

Thank you @appleboy2!

OpenStudy (appleboy2):

No problem!

OpenStudy (rebeccapink6398):

thx to you both @appleboy2 and @Bella5 :D..do you guys wanna know what I got on my test??

OpenStudy (appleboy2):

What? :D

OpenStudy (rebeccapink6398):

im in an online school..wanna see what we got right?

OpenStudy (appleboy2):

Ok!

OpenStudy (rebeccapink6398):

ok I got....

OpenStudy (rebeccapink6398):

..awe! 75%... I didn't pass it :( I have to get at least 80%

OpenStudy (appleboy2):

Oh no

OpenStudy (rebeccapink6398):

;(

OpenStudy (appleboy2):

How many questions were on the test?

OpenStudy (rebeccapink6398):

8

OpenStudy (appleboy2):

Can you review them?

OpenStudy (appleboy2):

Or see what you got wrong?

OpenStudy (rebeccapink6398):

yes one min..

OpenStudy (rebeccapink6398):

A pet show had 32 dogs, 24 cats, and 12 pets of other types. What was the ratio of dogs to all pets at the pet show? A. 4 : 9 B. 8 : 17 C. 8 : 9 D. 8 : 3 it was B

OpenStudy (rebeccapink6398):

Max bought a backpack that cost $40.00. The sales tax in his state is 8%. What is the final amount that Max has to pay for the backpack it was 43.20

OpenStudy (rebeccapink6398):

@appleboy2 u got ALL correct but those^^^^^^^^

OpenStudy (anonymous):

#8 was 74.20 ;(